# Competing charge density wave phases in YNiC2

### v1

@@ III.3 Specific heat and thermal expansion studies
Focusing on the CDW phase transitions, a close-up view on the specific heat anomalies presented in Fig. 4b reveals transition (i) from the orthorhombic parent-type to the $q_{1ic}$-type CDW state at $T_{1ic}=305$ K via a second order jump-like anomaly, $\Delta{C_{p}}\simeq 1.5(2)$ J mol-1K-1, and (ii) a first-order phase transition (FOT) from the $q_{1ic}$- to the $q_{2c}$-type CDW state at around $T_{2c}\simeq 272$ K is indicated by a peak-like anomaly. Considering the first order nature of the latter transition (ii) and associated latent heat, a standard evaluation of heat capacity data collected with a relaxation-type technique (as applied for Fig. 4b may yield erroneous results and, thus, an adapted slope analysis for heating and cooling curves across that specific heat anomaly has been applied as implemented for this specific purpose in the software of the PPMS heat capacity option). Specific heat data resulting from the slope analysis of one such heating-cooling cycle is displayed in Fig. 4c and reveals (within error bars imposed by thermal gradients) an essentially reversible nature of the nearly symmetric peak-shape anomaly when comparing the heating (dark blue) and cooling (light blue) data (results of additional cycles with varying initial temperatures are shown in Fig. S3b in SM [34]) with a very similar width ($\sim 266$ – 276 K) as indicated by the above discussed electrical resistivity data. The observed thermal reversibility across this transition is, within error bar, consistent with the above mentioned almost perfectly reversible behavior of the electrical resistivity, but nonetheless unexpected in context with a first order nature of that phase transition. A quantitative evaluation of the FOT anomaly in Fig. 4c yields associated enthalpy and entropy changes, $\Delta H\sim 30$ (3) J/mol and $\Delta S\sim 0.11$(1) J mol-1K-1, respectively.
Figure 5: (a) Temperature-dependent thermal expansion, $\Delta l/l(T)$, of YNiC2 along orthorhombic $a$-, $b$- and $c$-orientations. Solid lines are guides to the eye for revealing approximate slope changes across $T_{1ic}$. (b) Expanded view of the thermal expansion coefficients at the $q_{2c}$-type CDW ordering temperature $T_{2c}$.
@@
For both CDW transitions, the observed orientation dependent changes of the thermal expansion related to these phase transitions, when comparing signs of $\Delta l/l(T)$ anomalies along different orthorhombic orientations, imply reversed thermal shifts of the CDW phase transition upon stress or strain along these orientations (see below). At the second order CDW transition (i) at $T_{1ic}$, $\Delta l/l$ curves display rather subtle kink-like anomalies with an increase of slope upon increasing temperature for $a$- and $c$-axis data, but a decrease of slope for $b$-axis data. Experimental uncertainties caused by merging two different data sets right near transition (i) at $T_{1ic}$ impede a quantitative evaluation of the changes of thermal expansion coefficients $\Delta\alpha$. At the first order CDW transition (ii) at $T_{2c}$, jump-like anomalies $\Delta l_{a}/l_{a}\simeq+1.5\times 10^{-4}$ and $\Delta l_{b}/l_{b}\simeq+1.2\times 10^{-4}$ are observed for orthorhombic $a$- and $b$-orientations, i.e. an expansion when raising temperature across the FOT, but on the contrary a length contraction, $\Delta l_{c}/l_{c}\simeq-1.0\times 10^{-4}$, which add up to a relative volume expansion, $\Delta V/V\sim+1.7\times 10^{-4}$.
@@
Figure 6: (a) Electronic band structure and projected DOS in the orthorhombic type (a) parent state of YNiC2. The bands are colored with a linear combination of projections onto atomic orbitals. In particular, orange indicates the in plane combination, while blue indicates the out of plane combination, as defined in the text. (b) Sketch of the YNiC2 crystal structure highlighting the in plane (orange) and out of plane (blue) directions.
@@
A reversible behavior at a first order structure phase transition results from a coupling mechanism which stabilizes specific phase ratios within a finite temperature interval as proposed earlier for a (similar to present case) orthorhombic to monoclinic ferroelectric transition in a hafnium-zirconium oxide film [46]. Local compressive stress evolving from the orthorhombic to monoclinic structure transformation was suggested to stabilize an orthorhombic phase fraction across a wide temperature interval. The orthorhombic to monoclinic CDW phase transition of YNiC2, as detailed in the previous paragraph, is well suited to originate local strain patterns which, in course of energy minimization, stabilize a mixed pattern of $q_{1ic}$-type and $q_{2c}$-type CDW domains within a finite temperature interval, i.e. within the width of the transition. A simultaneous appearance of corresponding $q_{1ic}$ and $q_{2c}$ satellite reflections at around 275 K was also confirmed by single crystal diffraction data discussed in Sect. III.1.
Contrary to the single crystal case, aniosotropic thermal expansion of grains in a polycrystal will cause inter-grain strain already upon cooling within the orthorhombic parent state and, indeed, CDW phase transitions of polycrystalline YNiC2 extend over about twice as large temperature intervals (compare Ref. [39]). The nature of strain effects, originating from inter-grain strain as predetermined by thermal history in polycrystals versus energy minimizing orthorhombic to monoclinic phase coupling in monocrystals, is rather different and, thus, the suppression of hysteretic behavior expected to be less effective for polycrystals. The latter is supported by resistivity data of polycrystalline YNiC2 reported in Ref. [39] and by a reference heat capacity measurement of polycrystalline YNiC2 analogous to that in Fig. 4c, which however did not at all reproduce the essentially reversible behavior of the single crystal data (compare Figs. S3b and S3c in SM [34]).
@@ III.4 DFT study of electronic and vibrational properties
The choice of this unusual partitioning is motivated by the direct observation of the Kohn-Sham wavefunctions near the Fermi energy (Sect. IV of SM [34]), as well as the Wannier functions derived from those bands. The bands colored with these projections are shown in Fig. 6a; these are well separated in the in plane and out of plane characters, with the exception of a single band that changes character as it crosses the Fermi energy in the S–R and $\Gamma$–Z line. In addition to the bands, Fig. 6a shows the atom-projected density of states (DOS) where from -4 to -2 eV, the DOS exhibits mainly a Ni-$d$ character, while in the -2 to 2 eV window around the Fermi energy the atomic character is evenly spread between Y, Ni, and C, consistent with the hybrid character of these states.
@@
Figure 7: (a) Fermi surface decorated with an in plane projection, cut with a plane (dashed lines) (b) and (c) 2D cuts indicating nesting vectors with $q_{1}$ and $q_{2}$ momentum, respectively.
@@
Figure 8: Electronic band structure in $q_{1c}$-type (a) and $q_{2c}$-type (b) CDW state. The black line indicates the orthorhombic parent state, while the red line indicates the respective CDW phase.
@@
As we will discuss in the following, the change of band character is closely related to the CDW formation in YNiC2. The Fermi energy is crossed by bands in the S–R and $\Gamma$–Z lines, and in Z–T and T–Y (see Fig. 6a). Resulting FS sheets are shown in Fig. 7a with a color scheme indicating an in plane projection of the character of electronic states. There are two different sheets: a main FS sheet formed by open quasi-1D sheets which are connected by hourglass-like pillars, and a small electron pocket centered around the Z point. The orbital character of the main FS sheet is extremely uneven, ranging from mainly in plane around the $\Gamma$–Z direction (red color), to mainly out of plane around Z–T and T–Y (blue color). It was pointed out earlier, that the strong quasi-1D character of the FS creates proper conditions for nesting with two primary $q$-vectors (i.e. $q_{1}$ and $q_{2}$) which can be revealed via FS contours obtained at proper 2D cuts of the Brillouin zone [48, 49]. Accordingly, we show in Fig. 7b and  7c two slices of the Brillouin zone hosting $q_{1}$ and $q_{2}$ nesting vectors, which in our in plane projection connect states with intermediate (green-colored) character.
@@
In order to investigate the impact of CDW formation on the electronic structure of YNiC2, we directly compare the band structure of the orthorhombic parent phase with that of $q_{1c}$- and $q_{2c}$-type CDW phases in Fig. 8a and 8b, respectively, always following the paths in reciprocal space corresponding to the ones selected for the orthorhombic phase so that, folding aside, the bands lie on top of each other (see SM [34] Sect. II A for technical details). A common feature of both CDWs is that both open a gap in the $\Gamma$–Z and S–R lines, in correspondence to the region where the orbital character changes from in plane to out of plane. As a result, the only bands that previously changed character are now broken into two separate bands, each with its own well defined character. This separation of wave functions into bonding and anti-bonding states with distinctly distinguished in plane and out of plane character is highlighted in Fig. 9 via electron probability density plots as obtained from the Kohn-Sham wavefunction calculated for (a) the orthorhombic parent state and (b) the monoclinic $q_{2c}$-type CDW state (see also SM [34], Fig. S5 for projected band structures). This change of character is directly related to the direction of the atom displacements of the CDWs. Indeed, both are characterized mainly by the displacement of Ni in the $z$ direction, i.e. out of the $xy$ plane. This changes the hopping integrals, and lowers the energy of bands with out of plane character, and raises the in plane ones, which lowers the total electronic energy as compared to the undistorted phase. While both CDWs feature certain similarities, it is clearly noticeable that the electronic structure modification caused by the $q_{2c}$-type CDW is much more pronounced: 150 meV large gaps open in the S–R, T–Y, and $\Gamma$–Z lines, while for the $q_{1c}$-type CDW the main gap opening is visible slightly above Fermi, and is of around 25 meV. This leads to a depletion of states in an energy window around the Fermi energy (see Fig. S6 and S7 in SM [34]), which is larger in the $q_{2c}$-type CDW state.
@@
Figure 10 shows the Fermi surfaces decorated with a projection to in plane states for the orthorhombic parent (a), the $q_{1c}$- (b) and the $q_{2c}$-type (c) CDW states. In this perspective the separation of bands into in plane and out of plane character is even more evident, as CDW order splits the FS sheets into pockets of uniform character by removing a part of the quasi-planar sheets which connect them. This effect is already in action in the $q_{1c}$-type CDW, where a part of the quasi-1D sheet is already broken, and it further develops in the $q_{2c}$-type CDW state. Here, the FS becomes divided in a snake-like sheet connecting the pockets with in plane character (orange transparent box) and in the characteristic hourglass-shaped pillars with only out of plane character (azure transparent box), while an isolated pocket around the Z-point remains essentially unchanged.
@@
Figure 9: Comparison of electronic bands along $\Gamma$–S–R lines, colored with a linear combination of projections onto atomic orbitals, in the orthorhombic parent phase (a) and the $q_{2}$-type CDW phase (b). Aside, we show the probability density from the Kohn-Sham wavefunction at the wavevector and energy indicated with the black arrows.
@@
These results allow us to confirm that the tendency to form both CDW phases is inherent to the orthorhombic parent phase. In addition, phonon calculations for the two CDW states (see SM [34], Fig. S9) reveal that the precursors of the $q_{1c}$-type CDW are present in the $q_{2c}$-type one, and vice versa. Qualitatively, the tendency of the $q_{1ic}$-type CDW to form at higher temperature than the $q_{2c}$-type can be explained as a sort of intermediate phase which lowers the electronic energy less than the $q_{2c}$-type CDW phase, but with a higher electronic entropy gain.
@@
Figure 11: Phonon dispersions (a), decorated with electron-phonon coupling (red thick line), atom-projected phonon DOS, $F(\omega)$ (b), and electron-phonon spectral function, $\alpha^{2}F(\omega)$ (c), with integrated electron-phonon coupling. The projections onto Y, Ni, and C, are indicated as yellow, gray, and red shaded areas. (d) static limit of the bare susceptibility, $\zeta_{0}$, black line, and electron-phonon coupling for the first phonon mode, $\lambda(\mu=1)$, red line.
@@ IV Summary and discussion
The $q_{1}$-type CDW phase, appearing as the first CDW upon cooling from the orthorhombic parent state below $T_{1ic}$, weakly modifies the crystal structure producing a small incommensurate lattice modulation. The onset of this state opens a nearly isotropic electronic gap at Fermi energy. While a part of the high temperature FS is nested and removed, its overall character and topography is mostly conserved. The quasi-planar FS sheets running perpendicular to a of the undistorted orthorhombic state, responsible for favoring the $a$ direction in the terms of electrical resistivity, retain most of their area. For that reason, the $\rho_{b}$/$\rho_{a}$ and $\rho_{c}$/$\rho_{a}$ ratios, being the effective measure of transport anisotropy, remain rigid across the transition towards the $q_{1ic}$-type CDW state.
@@
An intriguing observation stems from the comparison of the electron-phonon coupling parameter, $\lambda$, calculated in the $q_{1c}$-approximate and the $q_{2c}$-type CDW states, respectively. Within each of the two CDW states, $\lambda(q)$ shows distinct maxima corresponding to the alternative CDW modulation. In particular, within the $q_{1c}$–approximate CDW state, $\zeta_{0}(q)$ shows a weakly developed feature at $q\simeq q_{2}$, while a well-defined, sharp peak is seen in $\lambda(q)$ at the same $q$-position. This observation suggests, that momentum dependent electron-phonon coupling plays a decisive role in the first-order incommensurate to commensurate CDW phase transition occurring at $T_{2c}$. Moreover, also in the $q_{2c}$-type CDW state, one can notice a clear maximum of $\lambda(q)$ at $q_{1}$. Interestingly, the signatures of such dormant adversary state precursors have already been observed in SmNiC2 and GdNiC2 compounds [21, 22]. In these materials, where only the $q_{1ic}$-type CDW develops a long range order, a weak X-ray diffuse scattering signal remains observable at the $q_{2c}$-position. Therefore, in YNiC2, and likely also in other intermediate members of the $R$NiC2 series, small details may decide on the result of the competition between $q_{1}$ and $q_{2}$ lattice modulations. The suppression of one of these entities by internal or external factors, such as rare-earth magnetism, uni-axial pressure or strain, may give the other one a potential to take over.

### v2

@@ III.3 Specific heat and thermal expansion studies
[…]
Figure 5: (a) Temperature-dependent thermal expansion, $\Delta l/l(T)$, of YNiC2 along orthorhombic $a$-, $b$- and $c$-orientations. Inset zooms on corresponding normalized thermal expansion data, $\Delta l(T)/\Delta l(380{\rm K})$, across the $q_{1ic}$-type CDW transition. Solid lines are guides to the eye for revealing slope changes across $T_{1ic}$. (b) Expanded view of the thermal expansion coefficients at the $q_{2c}$-type CDW ordering temperature $T_{2c}$.
@@
For both CDW transitions, the observed orientation dependent changes of the thermal expansion related to these phase transitions, when comparing signs of $\Delta l/l(T)$ anomalies along different orthorhombic orientations, imply reversed thermal shifts of the CDW phase transition upon stress or strain along these orientations (see below). At the second order CDW transition (i) at $T_{1ic}$, $\Delta l/l$ curves display rather subtle kink-like anomalies with an increase of slope upon increasing temperature for $a$- and $c$-axis data, but a decrease of slope for $b$-axis data (see inset in Fig. 5a). Experimental uncertainties caused by merging two different data sets right near transition (i) at $T_{1ic}$ impede a quantitative evaluation of the changes of thermal expansion coefficients $\Delta\alpha$. At the first order CDW transition (ii) at $T_{2c}$, jump-like anomalies $\Delta l_{a}/l_{a}\simeq+1.5\times 10^{-4}$ and $\Delta l_{b}/l_{b}\simeq+1.2\times 10^{-4}$ are observed for orthorhombic $a$- and $b$-orientations, i.e. an expansion when raising temperature across the FOT, but on the contrary a length contraction, $\Delta l_{c}/l_{c}\simeq-1.0\times 10^{-4}$, which add up to a relative volume expansion, $\Delta V/V\sim+1.7\times 10^{-4}$.
@@
Figure 6: (a) Electronic band structure and projected DOS in the orthorhombic type parent state of YNiC2. The bands are colored with a linear combination of projections onto atomic orbitals. In particular, orange indicates the Ni-C2 plane corresponding to the in plane combination of atomic orbitals, while blue indicates out of plane, as defined in the text. (b) Sketch of the YNiC2 crystal structure highlighting the in plane (orange) and out of plane (blue) areas.
@@
A reversible behavior in a first order structural phase transition results from a coupling mechanism which stabilizes specific phase ratios within a finite temperature interval as proposed earlier for a (similar to present case) orthorhombic to monoclinic ferroelectric transition in a hafnium-zirconium oxide film [46]. Local compressive stress evolving from the orthorhombic to monoclinic structure transformation was suggested to stabilize an orthorhombic phase fraction across a wide temperature interval. The orthorhombic to monoclinic CDW phase transition of YNiC2, as detailed in the previous paragraph, is well suited to originate local strain patterns which, in course of energy minimization, stabilize a mixed pattern of $q_{1ic}$-type and $q_{2c}$-type CDW domains within a finite temperature interval, i.e. within the width of the transition. A simultaneous appearance of corresponding $q_{1ic}$ and $q_{2c}$ satellite reflections at around 275 K was also confirmed by single crystal diffraction data discussed in Sect. III.1.
Contrary to the single crystal case, anisotropic thermal expansion of grains in a polycrystal will cause inter-grain strain already upon cooling within the orthorhombic parent state and, indeed, CDW phase transitions of polycrystalline YNiC2 extend over about twice as large temperature intervals (compare Ref. [39]). The nature of strain effects, originating from inter-grain strain as predetermined by thermal history in polycrystals versus energy minimizing orthorhombic to monoclinic phase coupling in monocrystals, is rather different and, thus, the suppression of hysteretic behavior expected to be less effective for polycrystals. The latter is supported by resistivity data of polycrystalline YNiC2 reported in Ref. [39] and by a reference heat capacity measurement of polycrystalline YNiC2 analogous to that in Fig. 4c, which however did not at all reproduce the essentially reversible behavior of the single crystal data (compare Figs. S3b and S3c in SM [34]).
@@ III.4 DFT study of electronic and vibrational properties
The choice of this unusual partitioning is motivated by the direct observation of the Kohn-Sham wavefunctions near the Fermi energy (Sect. IV of SM [34]), as well as the Wannier functions derived from those bands. The bands colored with these projections are shown in Fig. 6a; these are well separated in the in plane (i.e., in the Ni-C plane) and out of plane characters, with the exception of a single band that changes character as it crosses the Fermi energy in the S–R and $\Gamma$–Z line (coordinates of the reference points of the Brillouin zone are provided in Sect. II). In addition to the bands, Fig. 6a shows the atom-projected density of states (DOS) where from -4 to -2 eV, the DOS exhibits mainly a Ni-$d$ character, while in the -2 to 2 eV window around the Fermi energy the atomic character is evenly spread between Y, Ni, and C, consistent with the hybrid character of these states.
@@
Figure 7: (a) Fermi surface decorated with an in plane projection, cut with a plane (dashed lines) (b) and (c) 2D cuts indicating nesting vectors with $q_{\sim}(0.5,0.5,0)$ and $q_{2}\sim(0.5,0.5,0.5)$ momentum, respectively.
@@
Figure 8: Electronic band structure in $q_{1c}$-type (a) and $q_{2c}$-type (b) CDW state. The black line indicates the orthorhombic parent state, while the red line indicates the respective CDW phase. Blue arrows point at local band splittings.
@@
As discussed in the following, the change of band character is closely related to the CDW formation in YNiC2. The Fermi energy is crossed by bands in the S–R and $\Gamma$–Z lines, and in Z–T and T–Y (see Fig. 6a). Resulting FS sheets are shown in Fig. 7a with a color scheme indicating an in plane projection of the character of electronic states. There are two different sheets: a main FS sheet formed by open quasi-planar sheets which are connected by hourglass-like pillars, and a small electron pocket centered around the Z point. The orbital character of the main FS sheet is extremely uneven (see Fig. 7), ranging from mainly in plane around the $\Gamma$–Z direction (red color), to mainly out of plane around Z–T and T–Y (blue color). It was pointed out earlier, that the strong quasi-planar character of the FS creates proper conditions for nesting with two primary $q$-vectors (i.e. $q_{\sim}(0.5,0.5,0)$ and $q_{2}\sim(0.5,0.5,0.5)$) which can be revealed via FS contours obtained at proper 2D cuts of the Brillouin zone [48, 49]. Accordingly, we show in Fig. 7b and  7c two slices of the Brillouin zone hosting $q_{1}$ and $q_{2}$ nesting vectors, which in our in plane projection connect states with intermediate (green-colored) character.
In order to investigate the impact of CDW formation on the electronic structure of YNiC2, we directly compare the band structure of the orthorhombic parent phase with that of $q_{1c}$- and $q_{2c}$-type CDW phases in Fig. 8a and 8b, respectively, always following the paths in reciprocal space corresponding to the ones selected for the orthorhombic phase so that, folding aside, the bands lie on top of each other (see SM [34] Sect. II A for technical details, and Fig. S6b for the difference in DOS). Overall, both CDWs open a pseudogap (defined as a partial depletion of electronic states, where some portions of the Fermi surface develop an energy gap while others remain conductive [50]), and thus, system remains metallic. Here we define the CDW gap as the energy range over which the DOS in the distorted phase is lower than that of the undistorted parent phase (see Fig. S6). In momentum space, the band structure reveals local band splittings, e.g., along the T–Y (25 meV for q1c and 150 meV for q2c) and S–R lines, in correspondence to the regions where the orbital character changes from in plane to out of plane. As a result, the only bands that previously changed character are now broken into two separate bands, each with its own well-defined character. Although each individual splitting is small in magnitude, the cumulative effect of gap formation at different k-points (and energies) removes electronic states near the Fermi level and is seen as CDW pseudogap in the DOS.
@@
The separation of wave functions into bonding and anti-bonding states with distinctly distinguished in plane and out of plane character is highlighted in Fig. 9 via electron probability density plots as obtained from the Kohn-Sham wavefunction calculated for (a) the orthorhombic parent state and (b) the monoclinic $q_{2c}$-type CDW state (see also SM [34], Fig. S5 for projected band structures). This change of character is directly related to the direction of the atom displacements of the CDWs. Indeed, both are characterized mainly by the displacement of Ni in the $z$ direction, i.e. out of the $xy$ planes in which Ni and C atoms are located. This changes the hopping integrals, and lowers the energy of bands with out of plane character, and raises the in plane ones, which lowers the total electronic energy as compared to the undistorted phase. While both CDWs feature certain similarities, it is clearly noticeable that the electronic structure modification caused by the $q_{2c}$-type CDW is much more pronounced: more, and larger band splittings open in the S–R, T–Y, and $\Gamma$$Z$ lines mainly below the Fermi energy, while for the $q_{1c}$-type CDW the main splitting is visible slightly above the Fermi level. The resulting CDW pseudogap from the DOS analysis is about 283 and 414 meV, for $q_{1c}$- and $q_{2c}$-type CDW, respectively (see Fig. S6 in SM [34]).
@@
Figure 10 shows the Fermi surfaces decorated with a projection to in plane states for the orthorhombic parent (a), the $q_{1c}$- (b) and the $q_{2c}$-type (c) CDW states. In this perspective the separation of bands into in plane and out of plane character is even more evident, as CDW order splits the FS sheets into pockets of uniform character by removing a part of the quasi-planar sheets which connect them. This effect is already in action in the $q_{1c}$-type CDW, where a part of the quasi-planar sheet is already broken, and it further develops in the $q_{2c}$-type CDW state. Here, the FS becomes divided into a snake-like sheet connecting the pockets with in plane character (orange transparent box) and in the characteristic hourglass-shaped pillars with only out of plane character (azure transparent box), while an isolated pocket around the Z-point remains essentially unchanged.
@@
Figure 9: Comparison of electronic bands along $\Gamma$–S–R lines, colored with a linear combination of projections onto atomic orbitals, in the orthorhombic parent phase (a) and the $q_{2}$-type CDW phase (b). Aside, we show the probability density from the Kohn-Sham wavefunction indicated with the black arrows with momenta (0.00, 0.50, 0.13) and (-0.047, 0.38, 0.64) for the orthorhombic parent phase and $q_{2}$-type CDW phase, respectively.
@@
These results allow us to confirm that the tendency to form both CDW phases is inherent to the orthorhombic parent phase. In addition, phonon calculations for the two CDW states (see SM [34], Fig. S9) reveal that the precursors of the $q_{1c}$-type CDW are present in the $q_{2c}$-type one, and vice versa. Since all three phases are almost degenerate in energy (see Tab. III of SM [34]), and the depletion region of the electronic DOS in the $q_{1c}$-type CDW state is narrower than that of the $q_{2c}$-type CDW, the tendency of the $q_{1ic}$-type CDW to form at higher temperature than the $q_{2c}$-type can be qualitatively explained as a kind of intermediate phase with a slightly higher electronic entropy gain (see Fig. S7 of SM [34]).
Figure 11: Phonon dispersions (a), decorated with electron-phonon coupling (red thick line), atom-projected phonon DOS, $F(\omega)$ (b), and electron-phonon spectral function, $\alpha^{2}F(\omega)$ (c), with integrated electron-phonon coupling. The projections onto Y, Ni, and C, are indicated as turquise, gray, and red shaded areas. (d) static limit of the bare susceptibility, $\zeta_{0}$, black line, and electron-phonon coupling for the first phonon mode, $\lambda(\mu=1)$, red line.
@@ IV Summary and discussion
The $q_{1ic}$-type CDW phase, appearing as the first CDW upon cooling from the orthorhombic parent state below $T_{1ic}$, weakly modifies the crystal structure producing a small incommensurate lattice modulation. The onset of this state opens a nearly isotropic electronic gap at Fermi energy. While a part of the high temperature FS is nested and removed, its overall character and topography is mostly conserved. The quasi-planar FS sheets running perpendicular to a of the undistorted orthorhombic state, responsible for favoring the $a$ direction in the terms of electrical resistivity, retain most of their area. For that reason, the $\rho_{b}$/$\rho_{a}$ and $\rho_{c}$/$\rho_{a}$ ratios, being the effective measure of transport anisotropy, remain rigid across the transition towards the $q_{1ic}$-type CDW state.
@@
An intriguing observation stems from the comparison of the electron-phonon coupling parameter, $\lambda$, calculated in the $q_{1c}$-approximate and the $q_{2c}$-type CDW states, respectively. Within each of the two CDW states, $\lambda(q)$ shows distinct maxima corresponding to the alternative CDW modulation. In particular, within the $q_{1c}$–approximate CDW state, $\zeta_{0}(q)$ shows a weakly developed feature at $q\simeq q_{2}$, while a well-defined, sharp peak is seen in $\lambda(q)$ at the same $q$-position. This observation suggests that momentum dependent electron-phonon coupling plays a decisive role in the first-order incommensurate to commensurate CDW phase transition occurring at $T_{2c}$. Moreover, also in the $q_{2c}$-type CDW state, one can notice a clear maximum of $\lambda(q)$ at $q_{1}$. Interestingly, the signatures of such dormant adversary state precursors have already been observed in SmNiC2 and GdNiC2 compounds [21, 22]. In these materials, where only the $q_{1ic}$-type CDW develops a long range order, a weak X-ray diffuse scattering signal remains observable at the $q_{2c}$-position. Therefore, in YNiC2, and likely also in other intermediate members of the $R$NiC2 series, small details may decide on the result of the competition between $q_{1ic}$ and $q_{2c}$ lattice modulations. The suppression of one of these entities by internal or external factors, such as rare-earth magnetism, uni-axial pressure or strain, may give the other one a potential to take over.